OpenStudy (anonymous):

If F(x)=xcosx, what is the exact value of F′(0)?

OpenStudy (kagıtucak):

F'(x)=cosx-sinx.x F'(0)=1-0=0

OpenStudy (anonymous):

first differentiate it..... F(x)=xcosx then f'(x)=cosx-xsinx then substitute f'(0)

OpenStudy (goformit100):

F(x)=xcosx F'(x)=x(-sinx)+cosx = -x sinx + cos x F'(0)=-(0)( sin0) + cos 0 F'(0)= 0 + 1 = 1..............Answer

OpenStudy (goformit100):

@blueskies7 Check it Ma'am.

OpenStudy (anonymous):

thnaks!

OpenStudy (anonymous):

*thanks